What is the percentage composition of Na2SO4?





1
Expert's answer
2022-06-09T02:31:04-0400

Solution:

Mr(Na2SO4) = 2×Ar(Na) + Ar(S) + 4×Ar(O) = 2×22.99 + 32.065 + 4×15.999 = 142.041


%Na in Na2SO4 = [2×Ar(Na) / Mr(Na2SO4)] × 100%

%Na in Na2SO4 = [2×22.99 / 142.041] × 100% = 32.37%

%Na in Na2SO4 = 32.37%


%S in Na2SO4 = [Ar(S) / Mr(Na2SO4)] × 100%

%S in Na2SO4 = [32.065 / 142.041] × 100% = 22.57%

%S in Na2SO4 = 22.57%


%O in Na2SO4 = [4×Ar(O) / Mr(Na2SO4)] × 100%

%O in Na2SO4 = [4×15.999 / 142.041] × 100% = 45.06%

%O in Na2SO4 = 45.06%


The percentage composition of sodium sulfate (Na2SO4) is 32.37% sodium (Na), 22.57% sulfur (S), and 45.06% oxygen (O)